OpenStudy (anonymous):

The forces F 1, F 2, ..., F n acting at the same point P are said to be in equilibrium if the resultant force is zero, that is, if F 1 + F 2 + ... + F n = 0. F 1 = i + j, F 2 = i – j, F 3 = 9 i + 2 j Find the additional force required for the forces to be in equilibrium P.

OpenStudy (lastdaywork):

Let the required force be F = xi + yj Solve - 1 + 1 + 9 + x = 0 1 + (-1) + 2 + y = 0

OpenStudy (wolfe8):

I would, add them all up and see what is left, then multiply everything with -1.